# KARMA: Efficient Structural Defect Segmentation via Kolmogorov-Arnold Representation Learning

### v1

@@ Acknowledgments
A supplementary document with seven sections extends the main paper by presenting: (1) Theoretical analysis of KARMA including TiKAN convergence proofs, generalization bounds, and approximation error analysis (Section 1); (2) Detailed architectural design breakdown with low-rank factorization justification and computational complexity analysis (Section 2); (3) Complete experimental methodology covering statistical frameworks, ablation studies, and robustness testing (Section 3); (4) In-depth performance analysis of runtime, memory, and energy efficiency (Section 4), (5) Implementation guidelines for reproducibility and deployment optimization (Section 5); and (6) Extended results with comprehensive class-wise metrics, failure case analysis, and qualitative visualizations for CSDD and S2DS datasets (Sections 6–7).

### v2

@@ Acknowledgments
A supplementary document with seven sections extends the main paper by presenting: (1) Theoretical analysis of KARMA including TiKAN convergence proofs, generalization bounds, and approximation error analysis (Section 1); (2) Detailed architectural design breakdown with low-rank factorization justification and computational complexity analysis (Section 2); (3) Complete experimental methodology covering statistical frameworks, ablation studies, and robustness testing (Section 3); (4) In-depth performance analysis of runtime, memory, and energy efficiency (Section 4), (5) Implementation guidelines for reproducibility and deployment optimization (Section 5); and (6) Extended results with comprehensive class-wise metrics, failure case analysis and qualitative visualizations for CSDD and S2DS datasets (Sections 6–7).